An introduction to gp2c

By Bill Allombert and Ariel Pacetti

November 22, 2013

Contents

1 What is gp2c? 1 1.1 Installing gp2c . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 A gp2c tutorial 2 2.1 How can I compile and run my scripts? . . . . . . . . . . . . . . 2 2.2 How can I compile directly with gp2c? . . . . . . . . . . . . . . . 4 2.3 Using gp2c to find errors in GP scripts . . . . . . . . . . . . . . . 6 2.4 Using compiled functions in a new program . . . . . . . . . . . . 6 2.5 Hand-editing the C file generated by gp2c . . . . . . . . . . . . . 7

3 Advanced use of gp2c 7 3.1 gp2c types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 Type declaration . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.3 Effect of types declaration on default values . . . . . . . . . . . . 9 3.4 Type casting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.5 Example of optimisation . . . . . . . . . . . . . . . . . . . . . . . 9 3.6 Types and member functions . . . . . . . . . . . . . . . . . . . . 11

4 Common problems 11 4.1 Meta-commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2 Unsupported functions. . . . . . . . . . . . . . . . . . . . . . . . 12 4.3 Memory handling and global variables. . . . . . . . . . . . . . . . 12 4.4 GP lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.5 The install command . . . . . . . . . . . . . . . . . . . . . . . . . 13

5 Command-line options of gp2c 14

1 What is gp2c?

The gp2c compiler is a package for translating GP routines into the C programming language, so that they can be compiled and used with the PARI system or the GP calculator.

The main advantage of doing this is to speed up computations and to include your own routines within the preexisting GP ones. It may also find bugs in GP scripts.

This package (including the latest versions) can be obtained at the URL:

http://pari.math.u-bordeaux.fr/download.html#gp2c

1.1 Installing gp2c

After downloading the file gp2c-x.y.zplt.tar.gz (where x,y,z and t depend on the version), you first have to unzip the file with the command: gunzip gp2c-x.y.z plt.tar.gz

This will create the new file gp2c-x.y.zplt.tar. Next you have to extract the files with the tar program: tar -xvf gp2c-x.y.z plt.tar

Note: You can do both steps at once with GNU tar by using the command: tar -zxvf gp2c-x.y.z plt.tar.gz

This creates a directory gp2c-x.y.zplt , which contains the main gp2c files. Now you have to install the program.

You need the file pari.cfg. This file can be found in the PARI object directory and is installed in $prefix/lib/pari/.

Copy or link this file in the gp2c directory, be sure to call it pari.cfg. ln -s .../lib/pari/pari.cfg pari.cfg

Run ./configure, which will search for the PARI version and some other configuration tools of the system. To install the program, type make, and the program will be compiled. You can then run make check to verify that everything has gone fine (a bunch of OK’s should show up). All of this is completely standard, and you are now ready to use gp2c.

You can use gp2c directly from this directory or you can install it by running make install as root. If you do not install it, you can run it from the gp2c directory by typing ./gp2c

2 A gp2c tutorial

2.1 How can I compile and run my scripts?

The simplest way to use gp2c is to call gp2c-run. If you want to know what happens in detail, see next section.

To make the examples easier to follow, please move to the gp2c directory and link the root of your PARI source there: ln -s .../pari .

As an example, we will take the file pari/examples/squfof.gp, which is a simple implementation of the well-known SQUFOF factoring method of D. Shanks.

We just run the command: ./gp2c-run pari/examples/squfof.gp

After a little processing we get a GP session. But this session is special, because it contains the compiled squfof function. Hence we can do the following:

parisize = 4000000, primelimit = 500000
? squfof(3097180303181)
[419]
i = 596
Qfb(133225, 1719841, -261451, 0.E-28)
%1 = 1691693

Let’s try a bigger example:

? squfof(122294051504814979)
[20137]
  ***   the PARI stack overflows !
  current stack size: 4.0 Mbytes
  [hint] you can increase GP stack with allocatemem()
? allocatemem()
  ***   Warning: doubling stack size; new stack = 8.0 MBytes.
? squfof(122294051504814979)
[20137]
[20137, 3445]
i = 46474
Qfb(321233929, 131349818, -367273962, 0.E-28)
%2 = 73823023

We need a large stack because by default gp2c does not generate code to handle the stack (the so-called gerepile code). To instruct gp2c to add gerepile code automatically, we must use the -g option. So quit this GP session and launch a new one with -g. Oh well, before that type ls pari/examples/squfof.gp*

pari/examples/squfof.gp    pari/examples/squfof.gp.run
pari/examples/squfof.gp.c  pari/examples/squfof.gp.so
pari/examples/squfof.gp.o

These are the files generated by gp2c-run:

• pari/examples/squfof.gp.c is the C file generated by gp2c.

• pari/examples/squfof.gp.o is the object file generated by the C compiler.

• pari/examples/squfof.gp.so is the shared library generated by the linker.

• pari/examples/squfof.gp.run is a file that contains the commands needed to load the compiled functions inside GP.

It is the shared library which is used by GP.

Now let’s continue: ./gp2c-run -g pari/examples/squfof.gp

parisize = 4000000, primelimit = 500000
? squfof(122294051504814979)
[20137]
[20137, 3445]
i = 46474
Qfb(321233929, 131349818, -367273962, 0.E-28)
%1 = 73823023

This time it works with no difficulty using the default stack. We would like to know how much faster the compiled code runs, so we need to load the non compiled squfof file in GP:

? \r pari/examples/squfof.gp
  ***   unexpected character: squfof(n)=if(isprime(n),retur
                                       ^--------------------

Why?? Because squfof already exists as an installed function and GP refuses to overwrite it. To solve this problem, we will add a suffix to the name of the compiled function under GP. Quit the session and type: ./gp2c-run -g -s c pari/examples/squfof.gp

Now the function squfof is named squfof c instead, so we can do

parisize = 4000000, primelimit = 500000
? \r pari/examples/squfof.gp
? #
   timer = 1 (on)
? squfof(122294051504814979)
[20137]
[20137, 3445]
i = 46474
Qfb(321233929, 131349818, -367273962, 0.E-28)
time = 5,810 ms.
%1 = 73823023
? squfof_c(122294051504814979)
[20137]
[20137, 3445]
i = 46474
Qfb(321233929, 131349818, -367273962, 0.E-28)
time = 560 ms.
%2 = 73823023

So the compiled version is more than ten times faster than the noncompiled one. However for more complex examples, compiled code usually runs only three times faster on average.

An extra trick: once you have run gp2c-run on your script, it is compiled and you can use the compiled version outside gp2c-run in any GP session by loading the file with extension .gp.run. For example quit the gp2c-run session and start gp and do

parisize = 4000000, primelimit = 500000
? \r pari/examples/squfof.gp.run

Now you have access to the compiled function squfof c as well.

2.2 How can I compile directly with gp2c?

Now we want to compile directly with gp2c to understand what happens. We should run the command

./gp2c pari/examples/squfof.gp > squfof.gp.c

This creates a file squfof.gp.c in the gp2c directory. Now read this file with your favorite editor.

The first line is highly system-dependent, but should be similar to:

/*-*- compile-command: "/usr/bin/gcc -c -o pari/examples/squfof.gp.o
  -O3 -Wall -I/usr/local/include pari/examples/squfof.gp.c
  && /usr/bin/gcc -o pari/examples/squfof.gp.so
  -shared   pari/examples/squfof.gp.o"; -*-*/

This is the command needed to compile this C file to an object file with the C compiler and then to make a shared library with the linker. If you use emacs, typing ’M-x compile’ will know about this command, so you will just need to type Return to compile.

The second line is

#include <pari/pari.h>

This includes the PARI header files. It is important that the header files come from the same PARI version as GP, else it will create problems.

The next lines are

/*
GP;install("squfof","D0,G,p","squfof","./pari/examples/squfof.gp.so");
GP;install("init_squfof","v","init_squfof","./pari/.../squfof.gp.so");
*/

The characters ”GP;” denote a command that should be read by GP at start-up. Here, the install() commands above must be given to GP to let it know about functions defined by the library. gp2c-run copy such commands to

the file ./pari/examples/squfof.gp.run.

Please read the entry about the install() command in the PARI manual. The init squfof function is an initialization function that is created auto-

matically by gp2c to hold codes that is outside any function. Since in our case there are none, this is a dummy function. In other cases, it is essential. The next lines are

GEN squfof(GEN n, long prec);
void init_squfof(void);
/*End of prototype*/

This is the C prototypes of your functions. The rest of the file is the C code proper.

For teaching purpose, let’s run the command

./gp2c -s c pari/examples/squfof.gp > squfof2.gp.c

and look at the difference between squfof.gp.c and squfof2.gp.c: diff -u squfof.gp.c squfof2.gp.c

--- squfof.gp.c Tue Feb 26 13:44:42 2002
+++ squfof2.gp.c        Tue Feb 26 13:44:49 2002
@@ -1,8 +1,8 @@
 /*-*- compile-command: "/usr/bin/gcc -c -o pari/examples/squfof.gp.o
  -DMEMSTEP=1048576 -g -Wall -Wno-implicit  -I/usr/local/include
  pari/examples/squfof.gp.c && /usr/bin/ld -o pari/examples/squfof.gp.so
  -shared   pari/examples/squfof.gp.o"; -*-*/
 #include <pari/pari.h>
 /*
-GP;install("squfof","D0,G,p","squfof","./pari/examples/squfof.gp.so");
-GP;install("init_squfof","v","init_squfof","./pari/.../squfof.gp.so");
+GP;install("squfof","D0,G,p","squfof_c","./pari/...les/squfof.gp.so");
+GP;install("init_squfof","v","init_squfof_c","./par.../squfof.gp.so");
 */
 GEN squfof(GEN n, long prec);
 void init_squfof(void);

If you are not familiar with the diff utility, the above means that only the two lines starting with GP;install have changed. In fact squfof is still named squfof in the C file, but the install command tells GP to rename it squfof c in the GP session.

2.3 Using gp2c to find errors in GP scripts

The gp2c compiler can also be used to find errors in GP programs. For that we should use the -W option like in

./gp2c -W pari/examples/squfof.gp > squfof.gp.c

Warning:pari/examples/squfof.gp:7:variable undeclared
p
Warning:pari/examples/squfof.gp:11:variable undeclared
dd
Warning:pari/examples/squfof.gp:11:variable undeclared
d
Warning:pari/examples/squfof.gp:11:variable undeclared
b
...
Warning:pari/examples/squfof.gp:45:variable undeclared
b1

This option lists variables that are used but not declared. It is important to declare all your variables with local(), or with global(). For gp2c, undeclared variables are taken to be “formal variables” for polynomials. For example if you write a function to build a second degree polynomial like

                      pol(a,b,c)=a*x^2+b*x+c

you must not declare ’x’ here, since it stands for the formal variable x.

2.4 Using compiled functions in a new program

One you have successfully compiled and tested your functions you may want to reuse them in another GP program.

The best way is to copy the install commands of the functions you use at the start of the new program so that reading it will automatically load the compiled functions.

As an example, we write a simple program fact.gp that reads

install("squfof","D0,G,p","squfof","./pari/examples/squfof.gp.so");
fact_mersenne(p)=squfof(2^p-1)

and run GP:

parisize = 4000000, primelimit = 500000
? \rfact
? fact_mersenne(67)
i = 2418
Qfb(10825778209, 4021505768, -13258245519, 0.E-28)
%1 = 193707721

So all goes well. But what is even better is that gp2c understands the install command and will be able to compile this new program.

Also this particular example will fail because as stated above, PARI/GP already has a squfof function, and the linker will pick the wrong one, which is unfortunate.

So use the -p option to gp2c-run to change squfof to my squfof. ./gp2c-run -pmy -g pari/examples/squfof.gp

This option prefixes my to every GP name in the program so as to avoid name clashes. Change fact.gp to

install("my_squfof","D0,G,p","squfof","./pari/examples/squfof.gp.so");
fact_mersenne(p)=squfof(2^p-1)

and run ./gp2c-run -g fact.gp

parisize = 4000000, primelimit = 500000
? fact_mersenne(67)
i = 2418
Qfb(10825778209, 4021505768, -13258245519, 0.E-28)
%1 = 193707721

Nice isn’t it?

But it gets even better: instead of writing the install command directly in your script you can just load the squfof.gp.run using \r: just change fact.gp to

\r ./pari/examples/squfof.gp.run
fact_mersenne(p)=squfof(2^p-1)

2.5 Hand-editing the C file generated by gp2c

If you have some experience in PARI programming, you may want to manually edit the C file generated by gp2c, for example to improve memory handling. Here some tips:

• If you preserve the install() at the start of the file, you can use the command gp2c-run file.c to recompile your file and start a new GP session with your functions added, just as you use gp2c-run with GP scripts.

• More generally, gp2c-run automatically passes any line in the C file starting with ’GP;’ to GP at start-up.

• As explained in Section 2.2, under emacs you can type ’M-x compile’ to recompile the shared library.

3 Advanced use of gp2c

3.1 gp2c types

Internally gp2c assign types to objects. The most common types are given below:

name  description
void  like in C
bool  boolean, true (1) or false (0)
negbool antiboolean, true (0) or false (1)
small  C integer long
int  multiprecision integer
real  multiprecision floating point
mp  multiprecision number
var  variable
pol  polynomial
vecsmall vector of C long (t VECSMALL)
vec  vector and matrices (excluding vecsmall)
list  GP lists
str  characters string as a char *
genstr  characters string as a GEN (t STR)
gen  generic PARI object (GEN)
lg  length of object (returned by length)
typ  type of object (returned by type)

[figure: gen; mp; vec; pol; var; real; int; small; bool; negbool; void]

Table 1: Types preorder

Types are preordered as in Table 1. The complete preorder known by gp2c can be accessed by running gp2c -t.

Variables are typed. A variable can only take values having a type equal or lower than its type. By default, variables are of type gen.

3.2 Type declaration

To declare a variable as belonging to type type, use:

function(x:type,y:type=2) local(x:type, y:type=2) global(x:type, y:type=2) for(i:type=...

To declare several variables of the same type type at once, use:

local(x, y=2):type global(x, y=2):type You can even mix the two ways:

local(x, y:type2 =2):type1

will declare x to be of type type1 and y of type type2.

3.3 Effect of types declaration on default values

Under GP, all GP variables start with a default value, which is 0 for a local variable and ’v for a global variable v.

The gp2c compiler follows this rule for variables declared without a type. However, when a variable is declared with a type, gp2c will not assign it a default value. This means that the declaration local(g) is equivalent to local(g:gen=0), but not to local(g:gen), global(g) is equivalent to global(g:gen=’g), but not to global(g:gen), and f(g)=... is equivalent to f(g:gen=0)=..., but not to f(g:gen)=....

This rule was chosen for several reasons:

• The default value (0 or ’v) might not be an object suitable for the type in question. For example, local(v:vec) declares v as being of type vec. It would make no sense to initialize v to 0 since 0 does not belong to type vec. Similarly global(N:int) declares N as being of type int. It would make no sense to initialize N to ’N since ’N does not belong to type int.

• This allows defining GP functions with mandatory arguments. This way, GP will issue an error if a mandatory argument is missing. Without this rule, there is no way to tell apart 0 from a missing argument.

• This allows telling gp2c not to generate useless default values.

3.4 Type casting

Sometimes, we know a more precise type than the one the transtyping algorithm can derive. For example if x is a real number, its logarithm might be complex. However, if we are sure x is positive, the logarithm will be real.

To force an expression to belong to type type, use the syntax: expr:type gp2c will check types consistency and output warnings if necessary. For example f(x:int)=local(r:real); r=log(xˆ 2+1) gp2c will complain that the logarithm might not be real. Since xˆ 2+1 is always positive, we can write: f(x:int)=local(r:real); r=log(xˆ 2+1):real

3.5 Example of optimisation

Declaring the types of variables allow gp2c to perform some optimisations. For example, the following piece of GP code

rho(n)=
{
  local(x,y);

x=2; y=5;
while(gcd(y-x,n)==1,
  x=(x^2+1)%n;
  y=(y^2+1)%n; y=(y^2+1)%n
 );
gcd(n,y-x)
}

generates the following output:

GEN
rho(GEN n)
{
GEN x = gen_0, y = gen_0;
x = gen_2;
y = stoi(5);
while (gcmp1(ggcd(gsub(y, x), n)))
{
  x = gmod(gaddgs(gsqr(x), 1), n);
  y = gmod(gaddgs(gsqr(y), 1), n);
  y = gmod(gaddgs(gsqr(y), 1), n);
}
return ggcd(n, gsub(y, x));
}

The functions gsqr, gaddgs, gmod, ggcd are generic PARI functions that handle gen objects. Since we only want to factor integers with this method, we can declare n, x y of type int:

rho(n:int)=
{
local(x:int,y:int);
x=2; y=5;
while(gcd(y-x,n)==1,
  x=(x^2+1)%n;
  y=(y^2+1)%n; y=(y^2+1)%n
 );
gcd(n,y-x)
}

The new C code output by gp2c is:

GEN
rho(GEN n)        /* int */
{
GEN x, y;       /* int */
if (typ(n) != t_INT)
  pari_err(typeer, "rho");
x = gen_2;
y = stoi(5);

  while (gcmp1(gcdii(subii(y, x), n)))
  {
    x = modii(addis(sqri(x), 1), n);
    y = modii(addis(sqri(y), 1), n);
    y = modii(addis(sqri(y), 1), n);
  }
  return gcdii(n, subii(y, x));
}

Now, the code now uses the more specific functions sqri, addis, modii and gcdii.

The most efficient way to use typing is to declare some variables of type small. This way, these variables will be implemented by C long variables, which are faster than PARI integers and do not require garbage collecting. However, you will not be protected from integer overflow. For that reason, gp2c will automatically declare some loop indices of type small when the range cannot cause overflow. Sometimes gp2c can be too conservative but you can force a loop index to be small with the syntax for(i:small=a,b,...).

3.6 Types and member functions

For use with members functions, gp2c provides the following types:

nf for ordinary number fields, i.e., a result given by the GP function nfinit.

bnf for big number fields, i.e., a result given by the GP function bnfinit which includes class and unit group data.

bnr for ray class groups, i.e., a result given by the GP function bnrinit.

ell for elliptic curves, i.e., a result given by the GP function ellinit.

gal for galois extensions, i.e., a result given by the GP function galoisinit.

prid for prime ideals, i.e., a component of the result given by the GP function idealprimedec.

Members functions on typed objects are much more efficient.

4 Common problems

4.1 Meta-commands.

Meta-commands (commands starting with a \) other than \r are currently ignored by gp2c, though a warning will be issued, because it is not clear what they should do in a compiled program. Instead you probably want to run the meta-command in the GP session itself.

The meta-command \rinclude is replaced with the content of the file include (or include.gp) when gp2c reads the file. If you would prefer gp2c to link include.so to the program instead, see Section 2.4.

4.2 Unsupported functions.

The functions forell and forsubgroup are currently not implemented as an iterator but as a procedure with callbacks, which limits what you can do inside the loop.

Some functions are passed to GP by gp2c-run at start-up (using the GP; syntax) instead of being translated in C: install and addhelp. In practice, they can be considered as supported.

Some functions are built-in in GP, and so are not available to libpari programs. So if you use them gp2c generate a program that needs to run under GP, or with gp2c-run. Since the C prototype for those functions are not available, the C compiler will output a warning. This serves as a reminder of the problem. They are all the plotting functions and allocatemem, default, extern, input, quit, read, system, whatnow.

Some GP functions are not available for C programs, so the compiler cannot handle them. If you use them you will get the infamous ”unhandled letter in prototype” error. Sorry for the inconvenience. These are plot, ploth, plotrecth.

Also the functions read, eval, kill may compile fine but have a surprising behaviour in some case, because they may modify the state of the GP interpreter, not of the compiled program. Please see Section 4.3 for details. For example

f(n)=eval("n^2") is very different from f(n)=n^2.

The forstep function is supported when the step is a number. If it is a vector, you must add a tag :vec to make GP know about it like in

f(x)=
{
  local(v);
  v=[2,4,6,6,6,6,6,2,4,6,6]
  forstep(y=7,x,v:vec,print(y))
}

This is not needed if the step is a vector or a variable of type vec, but is needed if the step is only an expression which evaluates to a vector.

There is little that can be done about these problems without changing PARI/GP itself.

4.3 Memory handling and global variables.

While a lot of work has been done to ensure that gp2c handles global variables properly, the use of global variables is still a lot of trouble, so try to avoid them if you do not understand the implications on memory handling.

First, there is a common practice to use undeclared variables as formal variables, for example we assume x=’x and write a*x+b instead of a*’x+b. So gp2c will not raise an undeclared variable to the rank of global variable unless you declare it with the global() command, or you use it at toplevel (i.e. outside any function). See also Section 2.3

Second, global variables seen by a compiled function are C variables, not GP variables. There is no connection between the two. You may well have two variables with the same name and a different content. Currently GP knows only how to install functions, not variables, so you need to write compiled functions in order to access global variables under GP.

Basically, global variables are allocated in the main stack which is destroyed each time GP prints a new prompt. This means you must put all your commands on the same line. Also global variables must be initialized using the init <filename> function before being used, and are only supported with the -g flag.

So you end up doing gp2c-run -g global.gp

parisize = 4000000, primelimit = 500000
? init_global();myfunction(args);

Note that nothing prevents you from calling init global in the GP program. In that case, you can omit the parentheses (i.e, write init global, not init global()) so that you can still run your noncompiled program.

Another way to handle global variables is to use the clone function which copies a PARI object to the heap, hence avoids its destruction when GP prints a new prompt. You can use unclone to free a clone. Please read the PARI/GP manual for more information about clone.

A good use of clone is for initializing constant variables: for example in test/gp/initfunc.gp, the vector T is initialized by

T=clone([4,3,2,2,1,1,1,1,0,0,0,0,0,0,0,0])

You must still run the init <filename> after starting GP, but after that you can use T safely.

GP itself currently does not know about clone and unclone, but you can use dummy functions

clone(x)=x
unclone(x)=while(0,)

when running uncompiled.

4.4 GP lists

GP lists are not fully supported by gp2c. A partial support is available with the list type. You must tell gp2c that a variable will contain a list by using L:list inside a declaration, where L is the name of the variable as explained in Section 3.

Currently, assigning to a list element (L[2]=x) will not work and lists will not be freed unless you explicitly use listkill.

Note: The PARI user’s manual states that lists are useless in library mode.

4.5 The install command

The install command is interpreted as a gp2c directive. This allows using installed function in compiled programs, see Section 2.4.

However this has some side-effects:

• If present, the lib argument must be a string, not an expression that evaluate to a string.

• The install command is not compiled, instead it is added to the list of functions to install.

5 Command-line options of gp2c

Here is a brief description of the main options of gp2c, which can be seen with

./gp2c -h.

In Section 2.1 we saw how to use the -g option.

• -g tells gp2c to generate gerepile calls to clean up the PARI stack and reduce memory usage. You will probably need this option, although the C code will be easier to read or hand-edit without it.

• -ofile tells gp2c to write the generated C file to the file file instead of the standard output.

• -iN allows you to change the indentation of the generated C file. So if you want 4 spaces, just use the -i4 option with gp2c. The default is 2.

• -W is useful for debugging the .gp file, in the sense that it detects if some local variable is undeclared. For example, if the file algorithm.gp has a routine like

radical(x)=F=factor(x)[,1];prod(i=1,length(F),F[i])

The variable ’F’ is undeclared in this routine, so when running gp2c with the -W option it prints

• -L adds compiler directives to the code so that warning and error found by the compiler are prefixed with the line number in the original GP file instead of the C file.

• -C adds code to perform range checking for GP constructs like x[a] and x[a,b]. This also checks whether x has the correct type. By default gp2c does not perform such check, which can lead to a runtime crash with invalid code. This option causes a small runtime penalty and a large C code readability penalty.

Warning:algorithm.gp:1:variable undeclared F

At present, an undeclared variable is taken to be a ”formal variable” for polynomials by gp2c, so do not declare it if that is what you intend. For example in pol(a,b,c)=a*x^2+b*x+c you must not declare x since it stands for the formal variable ’x.

• -pprefix A problem with C is that it is subject to name clashes, i.e., if a GP variable in your routine has the same name as a C symbol in the pari library, the compiler will report strange errors. So this option changes ALL user variables and user routine names by adding a prefix prefix to them. For example the GP routine add(x,y) with -pmy will become the C function my add(x,y).

Try this option each time the compiler fails to compile gp2c output to see if there is a name conflict. If this is the case, change the name in your GP script. It may be difficult to find conflicting names if your compiler is not verbose enough and if you are not familiar with the PARI code and C in general.

Example of conflicting names are top,bot,prec,un, but there are thousands of others and they may be system-dependent.

• -ssuffix: Add suffix to the names of the installed functions under GP. This is to avoid clashes with the original GP script. For example, if you want to compare timings you may want to use the option -s c This does not affect the C code, only the install commands.

• -S: Assume strict prototypes for functions. This means that you have declared all the local variables with local and not as parameters. Also, when a parameter is not mandatory, a default value must be assigned to it in the declaration, like for ’flag’ in

test(data,flag=0)={CODE}

This does not affect the C code, only the install commands.

• -h gives the help.

• -v gives the gp2c version.

• -l prints a list of all the GP functions known by the compiler. So if a routine contains a GP routine not on this list, gp2c will show an error when trying to translate it. For example the routine forvec is not yet implemented, so you should not use it.

Reasons why a GP function may not be known by the compiler are:

– The function is not part of the PARI library. See Section 4.2

– You use the old PARI 1.39 function names instead of the new ones. gp2c currently does not know about the ’compat’ default. Use whatnow under GP to get the current name. For example, mod() is now Mod().

– You use a GP function that does not exists in the GP version gp2c was compiled against. Please recompile gp2c against this GP version. Normally no functions are added between two stable releases of GP with the same minor version number (say 2.1.1 and 2.1.2) so there is no need to recompile gp2c when you upgrade. But if you use the developement versions, you need to recompile. Also some new developement versions may break old versions of gp2c, so upgrade gp2c at the same time. However, if you want to compile scripts which do not use the new functions, you do not need to recompile. Note that you may use the GP environment variables to tell gp2c-run which GP to use.

– -t Output the table of types known to the compiler, see Section 3.